Determine the pH of a 0.161 M CH3CO2H M solution if the Ka of CH3CO2H is 1.8 x 10^-5
Answer #1

pH of weak acid (HA) like acetic acid

So, pH =1/2 log (1.08 x 10-5)1/2 log 0.161 = 2.483 + 0.396 = 2.88